Circular No. 8094

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003cb IN NGC 4885 Further to IAUC 8093, M. Schwartz and P. R. Holvorcem report the

LOTOSS discovery of a supernova (mag 17.9) on unfiltered Tenagra II CCD images taken on Mar. 13.4 and 15.4 UT. SN 2003cb is located at α = 13^{h}00^{m}33^{s}.25, δ = −6^{o}51′09′′.3 (equinox 2000.0), which is 10′′.1 west and 1′′.7 north of the center of NGC 4885. An image taken on Feb. 7.5 showed nothing at this position (limiting mag ∼ 19.3).

UNUSUAL VARIABLE OBJECT IN LEO Further to IAUC 8093, A. Becker also reports the discovery of an

unknown object, located at α = 10^{h}53^{m}45^{s}.79, δ = −5^{o}37′44′′.3 (equinox 2000.0), that shows unusual short-timescale variability and is possibly associated with a very faint red galaxy. Magnitudes for the object: Mar. 5.17 UT, V = [23.5; 5.19, V = 21.9; 5.20, B = 22.3; 5.21, B = 21.6; 5.22, B = 21.5; 5.22, B = 21.6; 5.23, B = 21.8; 7.3, V = [25; 10.4, V = [24.

COMET C/2002 V1 (NEAT) C. B. Cosmovici, Istituto di Fisica dello Spazio Interplanetario, Con-

siglio Nazionale delle Ricerche, Rome, reports that he and the rest of the ITASEL team of the Medicina 32-m radio telescope (M. Teodorani, S. Montebugnoli, and G. Maccaferri) detected the 22-GHz water-maser line in comet C/2002 V1 at perihelion, using a fast, high-resolution spectrum analyzer (8192 channels; bandwidth 8 MHz) on Feb. 17, 18, and 19 (when r = 0.101–0.121 AU), during the strong coronal mass ejection (CME) detected by the SOHO spacecraft. The beam size was 2′, corresponding to an observed region at the comet of ∼ 80000 km. The mean antenna temperature was 140 ± 38 mK, giving an integrated flux of 0.70 ± 0.17 Jy km/s (FHWM = 47 KHz). The neutral water molecules show a velocity shift of −12.2 km/s with respect to the nucleus, confirming the previously observed peculiar strong acceleration of neutral water molecules in C/1996 B2 at perihelion (Cosmovici et al. 1998, Planet. Space Sci. 46, 467; in C/1996 B2, the shift varied between 22 and 44 km/s). Neutral water molecules normally are ejected from the nucleus at velocities around 1 km/s. This peculiarity could be explained by assuming the sublimation of excited water molecules from accelerated ionized icy grains that form an extended source around the nucleus. The strong CME may have contributed to the acceleration mechanism. Evaluation of the water-production rate is in progress.

2003 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green